Based on past experiences, a bank believes that 10% of the people who receive loans will not make payments on time. The bank has recently approved 250 loans. What’s the probability that over 12% of these clients will not make timely payments? Give your answer to 4 decimal places.
Answer #1

Solution

\sigma\hat p =  \sqrt[p( 1 - p ) / n] = \sqrt [(0.10 * 0.90) / 250 ] = 0.019

P( \hat p > 0.12) = 1 - P( \hat p < 0.12)

= 1 - P(( \hat p - \mu \hat p ) / \sigma \hat p < (0.12 - 0.10) / 0.019)

= 1 - P(z < 1.054)

= 0.1459

Probability = 0.1459